### v1

# scSiameseClu: A Siamese Clustering Framework for Interpreting single-cell RNA Sequencing Data

Abstract
Single-cell RNA sequencing (scRNA-seq) reveals cell heterogeneity, with cell clustering playing a key role in identifying cell types and marker genes. Recent advances, especially graph neural networks (GNNs)-based methods, have significantly improved clustering performance. However, the analysis of scRNA-seq data remains challenging due to noise, sparsity, and high dimensionality. Compounding these challenges, GNNs often suffer from over-smoothing, limiting their ability to capture complex biological information. In response, we propose scSiameseClu, a novel Siamese Clustering framework for interpreting single-cell RNA-seq data, comprising of 3 key steps: (1) Dual Augmentation Module, which applies biologically informed perturbations to the gene expression matrix and cell graph relationships to enhance representation robustness; (2) Siamese Fusion Module, which combines cross-correlation refinement and adaptive information fusion to capture complex cellular relationships while mitigating over-smoothing; and (3) Optimal Transport Clustering, which utilizes Sinkhorn distance to efficiently align cluster assignments with predefined proportions while maintaining balance. Comprehensive evaluations on seven real-world datasets demonstrate that scSiameseClu outperforms state-of-the-art methods in single-cell clustering, cell type annotation, and cell type classification, providing a powerful tool for scRNA-seq data interpretation.
22footnotetext: The code is available at the link: https://github.com/XPgogogo/scSiameseClu.
(a) scNAME
(b) scGNN
Figure 1: Similarity distributions of cell embeddings learned by scNAME and scGNN on dataset Human liver cells.
1 Introduction
Single-cell RNA sequencing (scRNA-seq) technology represents a significant advancement in bioinformatics, enabling the capture of comprehensive genetic information from individual cells Shapiro et al. (2013); Yang et al. (2024). Cell clustering, a key step in single-cell RNA sequencing analysis, groups cells by their gene expression patterns to uncover the complex characteristics of distinct cell populations and provide insights into their biological functions and interactions Kiselev et al. (2019). Moreover, clustering gene expression patterns further contributes to various downstream tasks, such as marker gene identification and cell type annotation Haghverdi et al. (2016). Cluster analysis of scRNA-seq data has been a vibrant research area over the past decade.
In recent years, advanced computational methods have been increasingly explored to address the challenges of analyzing high-dimensional and sparse scRNA-seq data. Classical clustering methods, such as K-means, are straightforward and computationally efficient but struggle to capture the complex, nonlinear relationships in scRNA-seq data. Meanwhile, recent research has focused on applying deep learning frameworks Ning et al. (2024) for learning representations of scRNA-seq data and classifying samples into distinct clusters. A commonly used method is self-supervised learning methods, which uncover effective representations of scRNA-seq data by reconstructing the original input data Eraslan et al. (2019); Lopez et al. (2018); Tian et al. (2019). However, these methods primarily focus on extracting features from individual cells, overlooking the complex relationships between cells that are essential for understanding cellular diversity. Benefiting from the powerful utilization of graph information Ning et al. (2021), graph neural networks (GNNs) have been applied to analyze scRNA-seq data by modeling cells as nodes and their interactions as edges, effectively capturing both gene expressions and cell graphs to comprehensively represent cellular heterogeneity Luo et al. (2021); Wang et al. (2021b); Gan et al. (2022); Zhan et al. (2023).
Though good performance has been achieved, previous GNN-based works still face limitations when tackling the following challenges: (1) Deficient exploration on intercellular information: when applying GNN-based models to single-cell data, the graph construction process often overlooks the sparsity and noise inherent in scRNA-seq data, frequently leading to reduced model robustness. Most current methods rely on simple topologies, like cosine similarity, to build cell graphs from gene expression matrices Gan et al. (2022); Wang et al. (2021c), but do not handle the sparsity and noise inherent in scRNA-seq data. This leads to fragile graph representations, weakening GNNs’ ability to capture cellular interactions and limiting clustering performance and insights into cellular dynamics. (2) Insufficient prevention of representation collapse in cell embeddings: GNNs-based models for scRNA-seq data frequently encounter representation collapse, where embeddings of biologically distinct cells become overly similar in latent space. This issue arises from the inability of current methods to preserve the diversity of cell representations in sparse and noisy scRNA-seq data. As shown in Fig.1, both the deep learning-based scNAME Wan et al. (2022) and the GNN-based scGNN Wang et al. (2021b) exhibit varying degrees of representation collapse. The problem is more pronounced in scGNN, where the cosine similarity between almost all cell embeddings exceeds $0.5$, indicating a severe loss of diversity in cell embeddings. Such collapse diminishes the discriminative power of embeddings, blurring cluster boundaries and limiting the ability to distinguish different cell populations, ultimately degrading the effectiveness of GNN models in clustering tasks Zbontar et al. (2021).
To tackle the aforementioned challenge, we propose  scSiameseClu, a Siamese Clustering framework for interpreting single-cell RNA sequencing data. scSiameseClu is designed to capture and refine complex intercellular information while learning discriminative and robust representations across both gene and cellular features. scSiameseClu leverages three key components: dual augmentation module to enrich data, siamese fusion module to preserve critical information and reduce redundancy, and optimal transport clustering to align cluster distributions. It effectively explores intricate information, mitigates representation collapse, and achieves clearer cell population separation, excelling in scRNA-seq clustering and other biological tasks.
Our framework offers the following contributions:
  • We present scSiameseClu, a novel Siamese-based clustering framework tailored for scRNA-seq data that captures intricate information from gene expression and cell graphs to learn discriminative and robust cell embeddings, improving clustering outcomes and downstream tasks.
  • We introduce key components: (i) dual augmentation with distinct noise on gene expression and cell graphs to mitigate dropout effects and improve robustness; (ii) siamese fusion for cross-correlation and adaptive information fusion to enhance robustness; and (iii) optimal transport clustering to align distributions.
  • Experimental results on seven datasets demonstrate that scSiameseClu outperforms state-of-the-art (SOTA) methods in clustering and other biological tasks.
2 Related work
2.1 Deep Clustering for scRNA-seq.
In scRNA-seq data analysis, early clustering algorithms like Phenograph Levine et al. (2015), MAGIC Van Dijk et al. (2018), and Seurat Wang et al. (2025) utilize k-nearest neighbor (KNN) graphs to model cell relationships, while SIMLR Wang et al. (2018) and MPSSC Park and Zhao (2018) employ multiple kernel functions to derive robust similarity measures from different data representations. These methods, however, often struggle with high-dimensional data and complex nonlinear features, making them susceptible to noise. Deep learning approaches have since become prominent. DCA Eraslan et al. (2019) employs a Zero-Inflated Negative Binomial (ZINB) autoencoder to model scRNA-seq data distributions, capturing nonlinear gene dependencies. SCVI Lopez et al. (2018) and SCVIS Ding et al. (2018), which rely on autoencoders for dimensionality reduction, often face over-regularization due to their Gaussian distribution assumption. Additionally, scDeepCluster Tian et al. (2019) combines ZINB-based autoencoders with deep embedding clustering to optimize feature learning and clustering. Most approaches ignore inter-gene and inter-cell correlations, focusing solely on individual cell expression profiles.
2.2 Graph Clustering for scRNA-seq.
Recently, graph-based clustering methods have gained considerable attention for their effectiveness Tu et al. (2021); Liu et al. (2022); Ning et al. (2025). Methods like scGAE Luo et al. (2021) and graph-sc Ciortan and Defrance (2022) utilize graph autoencoders to embed scRNA-seq data while preserving topological structures. scGAC Cheng and Ma (2022) constructs cell graphs and employs a self-optimization approach for simultaneous representation learning and clustering. Notably, scGNN Wang et al. (2021b) uses GNNs and multi-modal autoencoders to aggregate cell-cell relationships and model gene expression patterns, while scDSC Gan et al. (2022) combines a ZINB model-based autoencoder with GNN modules using a mutually supervised strategy. These methods often struggle with representation collapse, limiting their ability to model cell relationships.
Figure 2: Overview architecture of scSiameseClu. It contains three components: (i) a data-augmented module, (ii) a Siamese fusion module, and (iii) an optimal transport clustering strategy for self-supervision learning.
3 Methodology
This section introduces the proposed scSiameseClu, including the highlighted dual augmentation module, siamese fusion module, optimal transport clustering. More details are provided in the supplementary material, e.g., gene expression autoencoders (GEAs) and cell graph autoencoders (CGAs).
3.1 Problem Formulation
Given a single-cell expression matrix $\mathbf{X}\in\mathbb{R}^{N\times D}$, where $x_{ij}\left(1\leq i\leq N,1\leq j\leq D\right)$ shows the expression of the $j$-th gene in the $i$-th cell. We construct an undirected graph $\mathcal{G_{C}}=(\mathcal{V},\mathcal{E})$ using KNN based on the gene expression matrix, where $\mathcal{V}=\{v_{1},v_{2},\cdots,v_{N}\}$ is the set of $N$ nodes, $\mathcal{E}$ is the set of edges. Each node corresponds to a single cell, and edges indicate similarity between cells, quantified with the Pearson Correlation Coefficient Ding et al. (2022); Lu et al. (2022). The KNN graph is defined with the preprocessed data matrix $\mathbf{\bar{X}}$, as specified in Section 4.1, and adjacency matrix $\mathbf{A}=(a_{ij})_{N\times N}$, where $a_{ij}=1$ if $v_{i}$ and $v_{j}$ are connected, and $a_{ij}=0$, otherwise. The associated degree matrix is $\mathbf{D}=diag(d_{1},d_{2},\cdots,d_{N})\in\mathbb{R}^{N\times N}$, where $d_{i}=\sum_{(v_{i},v_{j})\in\mathcal{E}}{a_{ij}}$ represents the degree of node $v_{i}$. To account for self-loops and normalize the adjacency matrix, we compute $\widetilde{\mathbf{A}}\in\mathbb{R}^{N\times N}$ as $\mathbf{D}^{-1}(\mathbf{A}+\mathbf{I})$, where $\mathbf{I}\in\mathbb{R}^{N\times N}$ is the identity matrix. Our goal is to partition the graph $\mathcal{G_{C}}$ into $\mathbf{C}$ categories by assigning each node $v_{i}\in\mathcal{V}$ to one of the $\mathbf{C}$ categories based on $\mathcal{E}$ and gene expression features.
3.2 Framework Overview.
We propose scSiameseClu, an enhanced graph autoencoder-based siamese clustering framework for interpreting scRNA-seq data. We aim to improve the performance of scRNA-seq clustering and related downstream tasks by learning more accurate and distinctive cell embeddings. As Fig. 2 shows, our framework consists of three components: (i) a dual augmentation module that enhances the gene expression matrix and the cell graph relationships; (ii) a siamese fusion module to map the augmented data into a lower-dimensional latent space, generating cell embeddings; and (iii) an optimal transport clustering strategy for self-supervision learning, which refines the latent representations and generates clustering assignments.
3.3 Dual Augmentation Module.
To enhance robustness against noise and generalization across diverse datasets, biologically plausible augmentations are introduced at both gene and cell levels. These augmentations capture the variation in scRNA-seq, helping the model extract biologically meaningful representations.
Gene Expression Augmentation. To enhance robustness at the gene level, biologically plausible augmentations are introduced by adding controlled noise to the gene expression profiles. Specifically, Gaussian noise is applied to simulate natural variability in gene expression. Let $\odot$ denote the Hadamard product Horn (1990), and $\mathbf{N}\in\mathbb{R}^{N\times D}$ represent a random noise matrix drawn from a Gaussian distribution $\mathcal{N}(1,0.1)$. The augmented gene expression matrix $\widetilde{\mathbf{X}}$ is computed as:
$\widetilde{\mathbf{X}}=\bar{\mathbf{X}}\odot\mathbf{N},$ (1)
as illustrated in Fig. 2, two augmented gene expression matrices, $\widetilde{\mathbf{X}}_{1}$ and $\widetilde{\mathbf{X}}_{2}$, were generated using Gaussian distributions with distinct parameters.
Cell Graph Augmentation. Cell graph augmentation aims to improve model robustness and enhance representation learning by introducing perturbations to the cell graph, specifically through edge perturbation and graph diffusion Wang et al. (2024). These two strategies provide distinct, yet complementary, perspectives on the cell graph, enabling the model to capture diverse interactions between cells.
(i) Edge perturbation. We employ an edge removal strategy to refine the cell graph while preserving the most meaningful and biologically relevant relationships. Removing weaker edges reduces graph noise and improves model robustness. In contrast, edge addition is avoided to prevent spurious connections that may distort the graph and compromise its biological validity. Specifically, a mask matrix $\mathbf{M}\in\mathbb{R}^{N\times N}$ is constructed based on the pairwise cosine similarity matrix computed in the latent space. The 10% of edges with the lowest similarity values are identified and removed using this mask. Finally, the adjacency matrix $\mathbf{A}^{r}\in\mathbb{R}^{N\times N}$ is normalized for proper scaling, calculated as:
$\mathbf{A}^{r}=\mathbf{D}^{-\frac{1}{2}}((\mathbf{A}\odot\mathbf{M})+\mathbf{I})\mathbf{D}^{-\frac{1}{2}}.$ (2)
(ii)Graph diffusion. We use a graph diffusion strategy to refine the cell graph by enhancing meaningful relationships between cells Hassani and Khasahmadi (2020). Specifically, the normalized adjacency matrix is transformed into a diffusion adjacency matrix using the Personalized PageRank (PPR) algorithm Page et al. (1998). The teleport probability controls information propagation across the graph, helping to emphasize biologically meaningful relationships. The diffusion adjacency matrix $\mathbf{A}^{d}$ would be computed as:
$\mathbf{A}^{d}=\alpha\left(\mathbf{I}-(1-\alpha)\left(\mathbf{D}^{-\frac{1}{2}}(\mathbf{A}+\mathbf{I})\mathbf{D}^{-\frac{1}{2}}\right)\right)^{-1},$ (3)
Thus, we define two distinct adjacency matrices, $\mathbf{A}^{r}$ and $\mathbf{A}^{d}$, from edge perturbation and graph diffusion, respectively.
3.4 Siamese Fusion Module.
We propose the Siamese Fusion Module (SFM), a novel framework that integrates cross-correlation refinement and adaptive information fusion to learn discriminative and robust sample representations and avoid the over-smoothing issue of GNN-based methods Ning et al. (2022). By minimizing the deviation of cross-correlation matrices from the ideal identity matrix across cells and genes and adaptive information fusion, SFM effectively reduces redundancy while preserving critical information Zbontar et al. (2021); Tu et al. (2021); Liu et al. (2022). This enhances the cell representations and avoids representation collapse, enabling more accurate cell population separation and improving clustering performance.
Cross-Correlation Refinement. The cross-correlation refinement process is designed to effectively encode and integrate the augmented data from genes and cells, ensuring the information from different augmentations is unified. We construct two gene expression autoencoders (GEAs) to process the augmented gene expression matrices $\widetilde{\mathbf{X}}_{1}$ and $\widetilde{\mathbf{X}}_{2}$, generating gene expression embeddings $\mathbf{Z}^{G1}\in\mathbb{R}^{N\times d}$ and $\mathbf{Z}^{G2}\in\mathbb{R}^{N\times d}$, where $d$ represents the latent embedding dimensionality. Similarly, we construct two cell graph autoencoders (CGAs) to process the augmented cell graph matrices ${\widetilde{\mathbf{X}}_{1},\mathbf{A}^{r}}$ and ${\widetilde{\mathbf{X}}_{2},\mathbf{A}^{d}}$, producing cell graph embeddings $\mathbf{Z}^{C1}\in\mathbb{R}^{N\times d}$ and $\mathbf{Z}^{C2}\in\mathbb{R}^{N\times d}$.
Cell Correlation Refinement. Cell correlation refinement (CCR) optimizes cell relationships across different augmented views by aligning corresponding embeddings and reducing redundant correlations. We can calculate the refined cell correlation matrix by $\mathbf{R}_{ij}^{C1}=\frac{\left({\mathbf{Z}}_{i}^{G1}\right)\left({\mathbf{Z}}_{j}^{G2}\right)^{\mathrm{T}}}{\left\|{\mathbf{Z}}_{i}^{G1}\right\|\left\|{\mathbf{Z}}_{j}^{G2}\right\|},\mathbf{R}_{ij}^{C2}=\frac{\left({\mathbf{Z}}_{i}^{C1}\right)\left({\mathbf{Z}}_{j}^{C2}\right)^{\mathrm{T}}}{\left\|{\mathbf{Z}}_{i}^{C1}\right\|\left\|{\mathbf{Z}}_{j}^{C2}\right\|},\forall i,j\in[1,N]$.
To ensure the cell correlation matrix $\mathbf{R}^{C1},\mathbf{R}^{C2}$ aligns with an identity matrix ${\mathbf{I^{C}}}\in\mathbb{R}^{N\times N}$, we minimize loss:
$\small\mathcal{L}_{Cor1}=\frac{1}{N^{2}}\sum\left(\mathbf{R}^{C1}-{\mathbf{I}}^{C}\right)^{2}+\frac{1}{N^{2}}\sum\left(\mathbf{R}^{C2}-{\mathbf{I}}^{C}\right)^{2}$ (4)
By ensuring the diagonal elements of $\mathbf{R}^{C1},\mathbf{R}^{C2}$ are equal to 1 and the off-diagonal elements are equal to 0, we can guarantee the alignment of embeddings for each view and minimize the consistency of embeddings for different cells across different views. This helps scSiameseClu reduce redundant information and learn more discriminative representations.
Latent Correlation Refinement. Similarly, latent correlation refinement (LCR) optimizes relationships between latent embeddings across different augmented views by aligning corresponding embeddings. First, we project gene expression embeddings $\mathbf{Z}^{G1}$ and $\mathbf{Z}^{G2}$ into $\widetilde{\mathbf{Z}}^{G1}$ and $\widetilde{\mathbf{Z}}^{G2}\in\mathbb{R}^{d\times K}$ with using a readout function $\mathcal{R}(\cdot):\mathbb{R}^{d\times N}\rightarrow\mathbb{R}^{d\times K}$ , formulated as $\widetilde{\mathbf{Z}}^{G1}=\mathcal{R}\left(\left(\mathbf{Z}^{G1}\right)^{\mathrm{T}}\right)$, $\widetilde{\mathbf{Z}}^{G1}=\mathcal{R}\left(\left(\mathbf{Z}^{G2}\right)^{\mathrm{T}}\right)$. Likewise, we can project cell graph embeddings $\mathbf{Z}^{C1}$ and $\mathbf{Z}^{C2}$ into $\widetilde{\mathbf{Z}}^{C1}=\mathcal{R}\left(\left(\mathbf{Z}^{C1}\right)^{\mathrm{T}}\right)$ and $\widetilde{\mathbf{Z}}^{C2}=\mathcal{R}\left(\left(\mathbf{Z}^{C2}\right)^{\mathrm{T}}\right)$.
Then we can calculate the refine latent correlation matrix by $\mathbf{R}_{ij}^{L1}=\frac{\left(\widetilde{\mathbf{Z}}_{i}^{G1}\right)\left(\widetilde{\mathbf{Z}}_{j}^{G2}\right)^{\mathrm{T}}}{\left\|\widetilde{\mathbf{Z}}_{i}^{G1}\right\|\left\|\widetilde{\mathbf{Z}}_{j}^{G2}\right\|},\mathbf{R}_{ij}^{L2}=\frac{\left(\widetilde{\mathbf{Z}}_{i}^{C1}\right)\left(\widetilde{\mathbf{Z}}_{j}^{C2}\right)^{\mathrm{T}}}{\left\|\widetilde{\mathbf{Z}}_{i}^{C1}\right\|\left\|\widetilde{\mathbf{Z}}_{j}^{C2}\right\|},\forall i,j\in[1,d]$, which denotes the correlation between the embeddings learned from enhanced gene expression matrix.
To ensure the latent correlation matrix $\mathbf{R}^{L1},\mathbf{R}^{L2}$ aligns with an identity matrix ${\mathbf{I}^{L}}\in\mathbb{R}^{d\times d}$, we minimize the loss:
$\small\mathcal{L}_{Cor2}=\frac{1}{d^{2}}\sum\left(\mathbf{R}^{L1}-{\mathbf{I}}^{L}\right)^{2}+\frac{1}{d^{2}}\sum\left(\mathbf{R}^{L2}-{\mathbf{I}}^{L}\right)^{2}$ (5)
Adaptive Information Fusion. To enhance clustering performance, we propose an adaptive information fusion mechanism that integrates cell relationships through embedding aggregation, self-correlation learning, and dynamic recombination. We first aggregate the embeddings achieved with GEA and CGA with a linear combination operation:
$\mathbf{Z}^{A}=\left(\mathbf{Z}^{G1}+\mathbf{Z}^{G2}+\mathbf{Z}^{C1}+\mathbf{Z}^{C2}\right)/4,$ (6)
Then, we apply a graph convolution-like operation (i.e., message passing) to process the combined information, enhancing the initial fused embedding $\mathbf{Z}^{A}\in\mathbb{R}^{N\times d}$. Specifically, we enhance $\mathbf{Z}^{A}$ by propagating cell graph $\widetilde{\mathbf{A}}$, i.e., $\mathbf{Z}^{E}=\widetilde{\mathbf{A}}\mathbf{Z}^{A}$. To capture the relationships among cell graphs in the enhanced embedding space, we compute a normalized self-correlation matrix $\mathbf{R^{S}}\in\mathbb{R}^{N\times N}$. Each element of $\mathbf{R^{S}}$ is defined as $\mathbf{R^{S}}_{ij}=\frac{exp{\left(\left(\mathbf{Z}^{E}{\left(\mathbf{Z}^{E}\right)}^{\mathbf{T}}\right)_{ij}\right)}}{\sum_{k=1}^{N}exp{\left(\left(\mathbf{Z}^{E}{\left(\mathbf{Z}^{E}\right)}^{\mathbf{T}}\right)_{ik}\right)}}$. The self-correlation matrix $\mathbf{R^{S}}$ encodes the relative similarity between cells, providing a mechanism to model relationships in the embedding space. Thus, we can recombine the embedding matrix $\mathbf{Z}^{E}$ to dynamically adjust the embedding representations. Specifically, the recombined embedding matrix $\mathbf{Z}^{R}$ is computed as $\mathbf{Z}^{R}=\mathbf{R^{S}}\mathbf{Z}^{E}$. Each cell integrates information from others through learned relationships, enhancing embeddings for clustering.
To preserve the initial fused embedding information while incorporating the recombined embeddings, we adopt a skip connection to fuse $\mathbf{Z}^{E}$ and $\mathbf{Z}^{R}$. The final clustering-oriented latent embeddings $\mathbf{Z}\in\mathbb{R}^{N\times d}$ are computed as:
$\mathbf{Z}=\mathbf{Z}^{E}+\beta\mathbf{Z}^{R},$ (7)
where $\beta$ is a learnable scale parameter, initialized to 0 and updated during training. This fusion mechanism effectively filters out redundant information and preserves discriminative features in the latent space, enabling the network to learn robust and meaningful representations that enhance clustering performance while avoiding representation collapse.
Propagated Regularization. The reconstruction loss, $\mathcal{L}_{REC}$ minimizes the joint mean square error (MSE) of reconstruction loss of gene expression and cell graph matrix. To further enhance representation quality, we introduce a propagation regularization. Drawing inspiration from Liu et al. (2022), formulated as follows:
$\mathcal{L}_{R}=JSD(\mathbf{Z},\widetilde{\mathbf{A}}\mathbf{Z}),$ (8)
where $JSD\left(\cdot\right)$ refers the Jensen-Shannon divergence Fuglede and Topsoe (2004). Utilizing Eq. 8, the network can capture long-range dependencies even with a shallow network architecture, thereby reducing over-smoothing as the propagated information deepens within the framework. Finally, the overall objective of the SFM module can be computed by:
$\mathcal{L}_{SFM}=\mathcal{L}_{Cor1}+\mathcal{L}_{Cor2}+\mathcal{L}_{REC}+\gamma\mathcal{L}_{R},$ (9)
where $\gamma$ is a trade-off parameter balancing the contribution of the propagation regularization.
3.5 Optimal Transport Clustering.
We utilize a self-supervised method to perform unsupervised clustering of scRNA-seq data Xie et al. (2016), which learn directly from the data without requiring labeled inputs. Specifically, we employ Student’s t-distribution Van der Maaten and Hinton (2008) as a kernel function to measure the similarity $q_{ij}$ between each cell embedding $h_{i}$ and the cluster center $c_{j}$. By assigning higher weights to closer points, this method effectively captures the non-linear relationships in scRNA-seq data. The assignment distribution is represented as a matrix $Q=\left[q_{ij}\right]$, where each element indicates the probability or similarity between sample $i$ and cluster center $j$. To further refine clustering, we compute a target distribution matrix $P=\left[p_{ij}\right]$, which sharpens the soft assignments in $Q$ by emphasizing high-confidence samples. This sharpening process improves cluster separation and ensures well-defined clusters. The clustering process is optimized by minimizing the divergence between $Q$ and $P$, iteratively refining results to group similar cells and separate dissimilar ones. In the target distribution $P$, each assignment in $Q$ is squared and normalized to enhance assignment confidence Bo et al. (2020).
To avoid degenerate solutions that would allocate all data points to a single label, we establish constraints that synchronize the label distribution with the mixing proportions. This ensures that each cell contributes equally to the loss calculation, improving clustering accuracy and preserving a balanced effect during learning process. We construct the target probability matrix $P$ using the following optimal transport strategy to ensure this alignment and enhance the robustness of our clustering outcomes:
$\min_{P}$ $-P*(\text{log}Q)$ (10)
s.t. $P\in\mathbb{R}_{+}^{N\times C},P\ \bm{1}_{C}=\bm{1}_{N}\ \text{ and }\ P^{T}\bm{1}_{N}=N\bm{\pi}.$
In this context, we regard the target distribution $P$ as the transport plan matrix derived from optimal transport theory, while $-\text{log}Q$ serves as the cost matrix. We impose the constraint $P^{T}\bm{1}_{N}=N\bm{\pi}$, where $\pi$ indicates the proportion of cells assigned to each cluster, estimated from intermediate clustering results. Thus, we can ensure that the resulting cluster distribution aligns with the defined mixing proportions. Given the substantial computational expense associated with direct optimization techniques, we utilize the Sinkhorn distance Sinkhorn (1967) to facilitate quicker optimization through an entropic constraint. The optimization problem, which includes a Lagrange multiplier for the entropy constraint is formulated as:
$\min_{P}$ $-P*(\text{log}Q)-\frac{1}{\lambda}\text{H}(P)$ (11)
s.t. $P\in\mathbb{R}_{+}^{N\times C},P\ \bm{1}_{C}=\bm{1}_{N}\ \text{ and }\ P^{T}\bm{1}_{N}=N\bm{\pi},$
where $\bm{H}$ is the entropy function measuring uncertainty in the distribution, and $\lambda$ is the smoothness parameter that maintains cluster balance. The transport plan $P$ is guaranteed to exist and be unique, with the solution efficiently obtained through Sinkhorn’s method Sinkhorn (1967) as follows:
$\hat{P}^{(t)}=\operatorname{diag}\left(\bm{u}^{(t)}\right)Q^{\lambda}\operatorname{diag}\left(\bm{v}^{(t)}\right)$ (12)
at each iteration $t$, $\bm{u}^{(t)}$ is updated as $\bm{1}_{N}/(Q^{\lambda}\bm{v}^{(t-1)})$ and $\bm{v}$ is calculated as $N_{\pi}/(Q^{\lambda}\bm{u}^{(t)})$. We begin with the initial value set as $\bm{v}^{(0)}=\bm{1}_{N}$. Upon convergence, the optimal transport plan matrix $\hat{P}$ is obtained.
During training, $\hat{P}$ is fixed and $Q$ is aligned with $\hat{P}$. This alignment is crucial for evaluating model performance, allowing us to define the clustering loss function as follows:
$\mathcal{L}_{OTC}=\text{KL}(\hat{P}\|Q)=\sum_{i}\sum_{j}\hat{p}_{ij}\log\frac{\hat{p}_{ij}}{q_{ij}}$ (13)
3.6 Joint Optimization.
The Zero-Inflated Negative Binomial (ZINB) loss $\mathcal{L}_{ZINB}$ is commonly used to handle the sparsity and overdispersion of scRNA-seq data by modeling excess zeros and variability, ensuring accurate data reconstruction Eraslan et al. (2019). Thus, in the proposed method, the overall optimization objective consists of three parts: the SFM loss, the OTC loss, and the ZINB loss, which are formulated as:
$\mathcal{L}=\mathcal{L}_{SFM}+\rho\mathcal{L}_{ZINB}+\sigma\mathcal{L}_{OTC},$ (14)
where the hyperparameters $\rho$ and $\sigma$ balance the components, facilitating effective embedding learning and clustering.
Dataset Metric pcaReduce SUSSC DEC scDeepCluster scDCC scNAME scDSC scGNN scCDCG OURS
Shekhar mouse retina cells ACC 21.27 $\pm$ 0.14 25.04 $\pm$ 0.73 24.55 $\pm$ 2.29 47.26 $\pm$ 3.34 74.14 $\pm$ 2.31 79.84 $\pm$ 2.77 67.54 $\pm$ 4.08 78.56 $\pm$ 3.95 76.04 $\pm$ 1.85 89.16 $\pm$ 3.52
NMI 25.54 $\pm$ 0.05 43.35 $\pm$ 0.36 40.32 $\pm$ 3.50 80.45 $\pm$ 1.28 81.20 $\pm$ 0.68 75.19 $\pm$ 11.89 62.46 $\pm$ 14.53 79.89 $\pm$ 5.28 76.71 $\pm$ 1.47 82.72 $\pm$ 1.22
ARI 10.01 $\pm$ 0.03 19.69 $\pm$ 0.52 17.96 $\pm$ 3.59 52.66 $\pm$ 5.77 62.44 $\pm$ 4.50 81.66 $\pm$ 7.05 54.53 $\pm$ 8.34 77.58 $\pm$ 9.59 60.65 $\pm$ 6.47 88.93 $\pm$ 2.15
Macosko mouse retina cells ACC 16.40 $\pm$ 0.33 30.34 $\pm$ 0.40 33.32 $\pm$ 2.97 54.45 $\pm$ 3.03 63.31 $\pm$ 11.09 78.97 $\pm$ 7.76 72.16 $\pm$ 0.77 62.80 $\pm$ 2.26 69.78 $\pm$ 0.70 87.13 $\pm$ 2.15
NMI 23.35 $\pm$ 0.19 50.72 $\pm$ 0.13 52.20 $\pm$ 2.14 76.83 $\pm$ 2.09 79.13 $\pm$ 0.94 82.64 $\pm$ 1.44 67.04 $\pm$ 0.49 68.25 $\pm$ 3.36 68.71 $\pm$ 2.52 86.62 $\pm$ 2.31
ARI 11.44 $\pm$ 0.79 18.89 $\pm$ 0.53 27.55 $\pm$ 6.65 40.90 $\pm$ 1.44 52.38 $\pm$ 2.78 84.72 $\pm$ 0.17 61.55 $\pm$ 0.82 74.01 $\pm$ 2.25 58.99 $\pm$ 2.69 90.92 $\pm$ 3.74
QS mouse lung cells ACC 33.70 $\pm$ 0.21 40.82 $\pm$ 0.10 38.07 $\pm$ 1.40 47.16 $\pm$ 3.35 48.29 $\pm$ 5.51 69.84 $\pm$ 2.77 67.54 $\pm$ 4.08 69.46 $\pm$ 1.88 70.58 $\pm$ 0.87 87.16 $\pm$ 1.52
NMI 16.10 $\pm$ 0.54 43.08 $\pm$ 0.01 37.27 $\pm$ 3.76 70.18 $\pm$ 2.24 50.40 $\pm$ 7.24 71.41 $\pm$ 5.40 69.48 $\pm$ 7.11 64.20 $\pm$ 2.10 70.13 $\pm$ 1.92 86.07 $\pm$ 1.04
ARI 11.68 $\pm$ 0.25 28.07 $\pm$ 0.04 23.08 $\pm$ 3.75 41.37 $\pm$ 4.07 42.76 $\pm$ 10.05 45.50 $\pm$ 3.20 76.76 $\pm$ 7.24 75.90 $\pm$ 0.97 80.67 $\pm$ 0.78 84. 49 $\pm$ 3.44
CITE-CMBC ACC 28.15 $\pm$ 0.07 47.19 $\pm$ 0.02 41.88 $\pm$ 5.35 67.59 $\pm$ 4.40 69.04 $\pm$ 3.46 60.22 $\pm$ 5.87 68.79 $\pm$ 0.87 66.71 $\pm$ 1.65 71.75 $\pm$ 1.8 74.70 $\pm$ 0.51
NMI 28.28 $\pm$ 0.19 60.14 $\pm$ 0.04 46.87 $\pm$ 9.23 73.16 $\pm$ 1.29 72.83 $\pm$ 1.31 48.98 $\pm$ 7.06 64.21 $\pm$ 3.25 61.70 $\pm$ 3.21 68.57 $\pm$ 1.7 68.71 $\pm$ 0.75
ARI 5.09 $\pm$ 0.04 32.84 $\pm$ 4.51 23.55 $\pm$ 10.25 51.21 $\pm$ 3.40 52.42 $\pm$ 2.60 66.06 $\pm$ 3.37 52.51 $\pm$ 2.15 66.25 $\pm$ 1.75 61.06 $\pm$ 1.4 67.56 $\pm$ 1.17
Human liver cells ACC 35.04 $\pm$ 0.14 47.96 $\pm$ 0.06 46.32 $\pm$ 6.76 61.32 $\pm$ 8.61 72.90 $\pm$ 4.02 74.55 $\pm$ 5.30 70.90 $\pm$ 2.23 72.33 $\pm$ 3.4 75.34 $\pm$ 1.7 88.33 $\pm$ 1.74
NMI 32.43 $\pm$ 0.33 67.17 $\pm$ 0.00 52.24 $\pm$ 3.44 78.09 $\pm$ 2.50 78.81 $\pm$ 1.50 77.42 $\pm$ 12.00 71.63 $\pm$ 2.75 73.56 $\pm$ 2.7 79.34 $\pm$ 2.6 88.82 $\pm$ 1.24
ARI 10.24 $\pm$ 1.42 35.27 $\pm$ 0.00 31.04 $\pm$ 10.90 58.28 $\pm$ 10.61 70.47 $\pm$ 1.93 79.91 $\pm$ 2.93 75.34 $\pm$ 3.56 76.01 $\pm$ 2.6 81.26 $\pm$ 2.7 91.90 $\pm$ 0.48
Human kidney cells ACC 41.38 $\pm$ 0.24 44.18 $\pm$ 0.22 38.91 $\pm$ 2.71 59.81 $\pm$ 4.69 67.89 $\pm$ 6.22 73.71 $\pm$ 3.90 75.32 $\pm$ 2.81 77.73 $\pm$ 1.41 79.55 $\pm$ 0.29 86.08 $\pm$ 0.89
NMI 24.31 $\pm$ 0.13 37.35 $\pm$ 0.02 28.78 $\pm$ 5.10 63.67 $\pm$ 3.70 66.53 $\pm$ 4.91 61.39 $\pm$ 4.70 69.41 $\pm$ 2.12 73.79 $\pm$ 1.07 67.98 $\pm$ 2.62 80.86 $\pm$ 0.96
ARI 24.56 $\pm$ 0.29 30.78 $\pm$ 0.02 19.97 $\pm$ 5.06 47.53 $\pm$ 5.44 50.37 $\pm$ 7.22 68.39 $\pm$ 2.41 67.74 $\pm$ 4.52 72.29 $\pm$ 1.59 64.91 $\pm$ 0.88 74.23 $\pm$ 2.07
Human pancreas cells ACC 51.17 $\pm$ 0.66 76.29 $\pm$ 0.02 72.22 $\pm$ 5.51 74.70 $\pm$ 2.69 86.65 $\pm$ 5.76 89.63 $\pm$ 5.83 79.42 $\pm$ 1.34 79.32 $\pm$ 3.96 92.65 $\pm$ 1.9 94.95 $\pm$ 0.15
NMI 57.71 $\pm$ 0.67 82.31 $\pm$ 0.13 76.29 $\pm$ 2.73 79.32 $\pm$ 0.32 83.81 $\pm$ 2.94 86.70 $\pm$ 7.79 75.89 $\pm$ 0.61 78.76 $\pm$ 5.60 85.81 $\pm$ 1.0 86.19 $\pm$ 0.34
ARI 37.00 $\pm$ 0.45 64.24 $\pm$ 0.01 61.76 $\pm$ 5.55 64.59 $\pm$ 2.49 79.83 $\pm$ 10.79 84.93 $\pm$ 3.13 75.42 $\pm$ 1.22 78.81 $\pm$ 3.17 91.37 $\pm$ 1.21 91.59 $\pm$ 0.35
Table 1: Clustering performance across seven datasets (mean $\pm$ standard deviation), with best results in bold and runner-up results in underline.
(a) scDeepCluster
(b) scNAME
(c) scGNN
(d) scCDCG
(e) scSiameseClu
Figure 3: Visualization of scSiameseClu and four typical baselines on human liver cells in 2D t-SNE projection. Each point represents a cell, while each color represents a predicted cell type.
(a) Similarity distribution
(b) Similarity heatmap
Figure 4: Distribution plot and heat map of cell similarities in latent space learned by scSiameseClu on the  Human liver cells dataset.
4 Experiments
In this section, we first validate our method through extensive experiments, demonstrating its superior performance in scRNA-seq clustering. We then show that the embeddings generated by our method effectively mitigate the embedding collapse issue in graph-based models. We evaluate its utility in downstream tasks, like cell type annotation, and perform ablation studies to assess the contributions of individual components. Additional details are in the supplementary material.
4.1 Experimental Setup.
Dataset Preprocessing. We evaluated scSiameseClu on seven real scRNA-seq datasets. First, genes expressed in fewer than three cells were filtered out, followed by normalization, log-transformed (logTPM), and selection of highly variable genes based on predefined mean and dispersion thresholds. Finally, the preprocessed data were used as input.
Baseline Methods. To verify the superior performance of the proposed approach, we compare it against nine competing SOTA clustering methods. Among them, pcaReduce Žurauskienė and Yau (2016) and SUSSC Wang et al. (2021a) represent early traditional clustering techniques. Additionally, we consider four deep neural network-based clustering methods, DEC Xie et al. (2016), scDeepCluster Tian et al. (2019), scDCC Tian et al. (2021), and scNAME Wan et al. (2022), which train an autoencoder to obtain representations followed by clustering. scDSC Gan et al. (2022), scGNN Wang et al. (2021b), and scCDCG Xu et al. (2024) are representative GNN-based clustering methods that incorporate both node attributes and structural information for latent representation.
Implementation Details. We implemented  scSiameseClu in Python 3.7 using PyTorch, with DCRN Liu et al. (2022) as the backbone. Each experiment was repeated 10 times, reporting the mean and variance. Detailed parameter settings are provided in the supplementary material.
Evaluation Metrics. To illustrate the effectiveness of our approach, we evaluate clustering performance through three widely recognized metrics: Accuracy (ACC), Normalized Mutual Information (NMI) Strehl and Ghosh (2002), and Adjusted Rand Index (ARI) Vinh et al. (2009). Higher values for these metrics indicate better clustering results.
(a) scDeepCluster
(b) scNAME
(c) scGNN
(d) scCDCG
(e) scSiameseClu
Figure 5: Cell type annotation: overlap of top 50 DEGs in clusters versus gold standard cell types (similarity = overlapping DEGs/50).
(a) Gold Standard
(b) scSiameseClu
Figure 6: Top 3 DEGs per cell type annotated by scSiameseClu and gold standard.
4.2 Overall Performance.
Quantitative Analysis. Tab. 1 compares the clustering performance of our method with nine competitive baselines across seven benchmark datasets. Based on the results, we can observe that: 1) scSiameseClu significantly outperforms traditional single-cell clustering methods, demonstrating its superior ability to handle scRNA-seq data. 2) Compared to deep learning-based methods, which rely solely on node information for clustering, scSiameseClu achieves consistently better results, highlighting its ability to leverage both gene and cell information effectively. 3) While GNN-based methods incorporate cell graph information, they suffer from redundant embeddings and representation collapse. In contrast, scSiameseClu learns more discriminative representations of scRNA-seq data, thereby enhancing clustering performance. Overall, scSiameseClu  outperforms the competing methods on three metrics across all datasets. On average, scSiameseClu achieves a significant average improvement of 8.41% in ACC, 5.89% in NMI, 4.34% in ARI than the second-best method.
Qualitative Analysis. To evaluate clustering performance, we used t-SNE to project the learned embeddings from different methods into a two-dimensional space on the Human liver cells dataset. As Fig. 3 shows, scSiameseClu produces well-separated clusters with clear boundaries, effectively distinguishing different cell types regardless of cluster size. In contrast, methods like scDeepCluster show dispersed clusters, failing to group similar cell types. We further analyze the embeddings from  scSiameseClu with the cosine similarity matrices. As shown in Fig. 4, the bimodal distribution of cosine similarities (approaching a multivariate Gaussian) and the distinct heatmap patterns preserve meaningful cluster structures in the latent space.  scSiameseClu mitigates representation collapse and effectively captures latent cluster structures, unlike other deep learning and GNN-based methods (Fig1).
4.3 Biological Analysis.
Cell Type Annotation. To demonstrate that our clustering method enables more effective downstream analysis, we first identified differentially expressed genes (DEGs) and marker genes for each cluster using the "FindAllMarkers" function in Seurat Butler et al. (2018). Then, we compared the top 50 marker genes identified by scSiameseClu and other methods to the gold standard on the human pancreas cells dataset, and our method accurately annotated clusters with cell types, achieving over 90% similarity in marker genes for most clusters. As Fig. 5 shows, cluster 1 was identified as endothelial cell type, and cluster 2 as type B pancreatic cells, etc. Additionally, Fig. 6 shows scSiameseClu identifies marker genes for each cluster, e.g., cluster 6 is characterized as pancreatic A cell type with marker genes TTR, GCG, and IRX2. Our method effectively matched clusters with known cell types, providing reliable support for mechanistic studies.
Metric pcaReduce scDCC scDSC scCDCG scGNN scSiameseClu
Accuracy 83.17 86.34 91.71 98.05 99.02 99.51
Precision 57.40 66.69 89.30 92.70 88.50 99.69
Recall 49.34 67.65 79.41 93.45 98.98 99.09
F1 48.17 65.82 78.84 93.06 99.23 99.36
Table 2: Classification performance comparison.
Cell Type Classification. To evaluate the discriminative power and generalization ability of the learned representations, we conducted classification experiments. Tab. 2 shows that scSiameseClu outperforms the baseline models on human pancreatic cells dataset across four metrics, including accuracy and F1-score. This demonstrates  scSiameseClu’s superior performance in identifying cellular heterogeneity and the robustness and effectiveness in cell type classification.
4.4 Ablation Study
Framework Component Study. To evaluate the effectiveness of each component in  scSiameseClu, we perform an ablation study on  Shekhar mouse retina cells dataset. Specifically, we denote the model without SFM loss as scSiameseClu-w/o SFM, while scSiameseClu-w/o ZINB and scSiameseClu-w/o OTC denote the model without ZINB loss and OTC loss, respectively. Likewise, scSiameseClu refers to the complete model with all components included. According to Fig. 7(a), each component significantly enhances the model’s performance, demonstrating their importance in improving efficiency and outperforming the baseline.
SFM Module Evaluation. As shown in Fig.7(a), SFM module plays a significant role in improving overall performance. To further validate the importance of its components, we compare our method (scSiameseClu) with its baseline through ablation experiments (Fig. 7(b)). Specifically, w/o SFM-C and w/o SFM-L denote the model without the cell correlation and latent correlation, respectively. Likewise, we denote the model without propagated regularization as w/o SFM-R, without MSE reconstruction loss as w/o SFM-REC. scSiameseClu also all these components. Fig. 7(b) shows that each SFM component improves performance, indicating that that scSiameseClu effectively integrates both gene and cell information.
(a) Major components evaluaion
(b) Component Effects in SFM
Figure 7: Ablation study on  Shekhar mouse retina cells dataset.
5 Conclusion
scSiameseClu integrates dual augmentation module, siamese fusion module, and optimal transport clustering to enhance representation learning while preserving biological relevance. Experimental results on seven datasets demonstrate that scSiameseClu not only outperforms nine baselines on scRNA-seq clustering but also alleviates the representation collapse issue common in GNN-based approaches. In addition, we conduct biological analyses, including cell type annotation and classification, underscoring the  scSiameseClu’s potential as a powerful tool for advancing single-cell transcriptomics.
6 Acknowledgements
This work is partially supported by the National Natural Science Foundation of China (Grant Nos. 92470204 and 62406306), the National Key Research and Development Program of China (Grant No. 2024YFF0729201), and the Science and Technology Development Fund (FDCT), Macau SAR (file no. 0123/2023/RIA2, 001/2024/SKL).
Supplementary Materials for scSiameseClu
Appendix A Data Details
A.1 Characteristic of scRNA-seq Data
Data generated from large-scale parallel scRNA-seq experiments forms a high-dimensional, high-sparsity gene expression matrix, where rows represent cells and columns correspond to genes. Most cell information is embedded in the non-zero expression values, with nearly 90% of the data being zero. These datasets are further characterized by high dropout rates, substantial noise, and complex structures, complicating the process of representation learning and necessitating advanced modeling techniques.
Technical biases inherent in transcriptome sequencing amplify data noise, leading to interference in the extraction of meaningful biological signals. The sparsity of the data not only introduces computational challenges, but also blurs cellular boundaries, reducing the precision of downstream clustering and annotation tasks. In addition, the growing coverage of genes increases the complexity of intercellular relationships, making it more difficult to accurately capture subtle interactions and similarities between cells.
In this context, methods that effectively capture intercellular relationships, minimize noise interference, and enhance the interpretability of scRNA-seq data are critical for improving clustering outcomes and enabling reliable biological insights.
A.2 Experimental Dataset
We evaluated the performance of the proposed scSiameseClu on seven real scRNA-seq datasets by conducting extensive experiments. Tab. 3 summarizes the brief information of these real datasets. These datasets originate from various sequencing methods. Of the seven datasets, three are derived from mouse samples and four from human samples. These datasets encompass a diverse array of cell types, such as retina, lung, liver, kidney, pancreas, and peripheral blood mononuclear cells.
Datasets Sequencing method Sample size No. of genes No. of groups Sparsity rate(%)
Shekhar mouse retina cells Shekhar et al. (2016) Drop-seq 27499 13166 19 93.33
Macosko mouse retina cells Macosko et al. (2015) Drop-seq 14653 11422 39 88.34
QS mouse lung cells Consortium et al. (2018) Smart-seq2 1676 23341 11 89.08
CITE-CMBC Zheng et al. (2017) 10X genomics 8617 2000 15 93.26
Human liver cells MacParland et al. (2018) 10X genomics 8444 4999 11 90.77
Human kidney cells Young et al. (2018) 10X genomics 5685 25125 11 92.92
Human pancreas cells Muraro et al. (2016) CEL-seq2 2122 19046 9 73.02
Table 3: Summary of the scRNA-seq datasets
Appendix B Abbreviations
Tab. 4 lists all abbreviations.
Abbreviation Full Form
scRNA-seq single-cell RNA sequencing
ZINB zero-inflated negative binomial
GNN graph neural network
SOTA state-of-the-art
KL Kullback-Leibler
ACC accuracy
NMI normalized mutual information
ARI adjusted rand index
GEA gene expression autoencoders
CGA cell graph autoencoders
CCR cell correlation refinement
LCR latent correlation refinement
Table 4: List of Abbreviations
Appendix C Method Details
C.1 Notation Summary
Tab. 5 summarizes the notations of all characters.
Notations Meaning
$\mathbf{X}\in\mathbb{R}^{N\times D}$ Preprocessed matrix
$\mathbf{A}\in\mathbb{R}^{N\times N}$ Adjacency matrix (Cell graph)
$\mathbf{I}\in\mathbb{R}^{N\times N}$ Identity matrix
$\mathbf{D}\in\mathbb{R}^{N\times N}$ Degree matrix
$\widetilde{\mathbf{A}}\in\mathbb{R}^{N\times N}$ Normalized adjacency matrix
$\mathbf{A}^{r}\in\mathbb{R}^{N\times N}$ Edge-removed adjacency matrix
$\mathbf{A}^{d}\in\mathbb{R}^{N\times N}$ Graph diffusion adjacency matrix
$\widehat{\mathbf{X}}\in\mathbb{R}^{N\times D}$ Reconstructed attribute matrix
$\widehat{\mathbf{A}}\in\mathbb{R}^{N\times N}$ Reconstructed adjacency matrix
$\mathbf{Z}\in\mathbb{R}^{N\times d}$ Clustering-oriented latent embedding
$\mathbf{Z}^{C}\in\mathbb{R}^{N\times d}$ Latent embedding of GEA
$\mathbf{Z}^{G}\in\mathbb{R}^{N\times d}$ Latent embedding of CGA
$\mathbf{R}^{\mathcal{C}}\in\mathbb{R}^{N\times N}$ Refined cell correlation matrix
$\mathbf{R}^{\mathcal{G}}\in\mathbb{R}^{d\times d}$ Refined latent correlation matrix
$\mathbf{Q}\in\mathbb{R}^{N\times C}$ Soft assignment distribution
$\mathbf{P}\in\mathbb{R}^{N\times C}$ Target distribution
Table 5: Notation summary
C.2 Data Augmentation
Tab. 6 shows the overview of data augmentations in this work.
Data augmentation Type Operation Object
Feature perturbation Feature-space Nodes
Edge perturbation structure-space Edges
Graph diffusion structure-space Edges
Table 6: Overview of data augmentations.
C.3 Gene expression and Cell Graph Autoencoders
As shown in the framework diagram of scSiameseClu, the encoder part adopts a Siamese network structure. With parameter sharing, each branch is equipped with gene expression antoencoders (GEAs) and cell graph autoencoders (CGAs) to respectively extract genes and cells information. In the decoder part, we use clustering-oriented latent embedding $\mathbf{Z}$ and normalized adjacency matrix $\widetilde{\mathbf{A}}$ as the input and reconstruct the inputs of the two sub-networks through the decoders of GEA and CGA, respectively. Unlike conventional autoencoders, GEA’s decoder has three output layers, modeling highly sparse and overdispersed scRNA-seq data with a ZINB-based distribution. Furthermore, CGA utilizes a symmetric architecture for reconstructing both the weighted attribute matrix and the adjacency matrix. This approach enhances CGA’s ability to capture both structural and attribute features of the graph, leading to more precise and comprehensive information reconstruction and a substantial enhancement in the network’s generalization capability.
Gene Expression Autoencoders. Gene expression autoencoders are proposed to learn the representative embedding of scRNA-seq expression through stacked three layers of dense networks in both the encoder and decoder. Taking perturbated attribute matrix $\widetilde{\mathbf{X}}$ as inputs, we denote the feature encoder and decoder functions as:
$\mathbf{Z}_{G}=f_{enc}\left(W_{GEA}\mathbf{X}+b\right)$ (15)
$\mathbf{H}=f_{dec}\left(W_{GEA}^{{}^{\prime}}\mathbf{Z}_{G}+b^{{}^{\prime}}\right),$
where $b$ and $b^{{}^{\prime}}$, respectively, represent the bias vectors of the encoder and the decoder, $W_{GEA}$ and $W_{GEA}^{{}^{\prime}}$ are the weights of layers learned from network training.
Cell Graph Autoencoder. Cell graph autoencoders are designed to model complex relationships and interactions among cells, using perturbated attribute matrix $\widetilde{\mathbf{X}}$ and the edge perturbation adjacency matrix $\mathbf{A}^{r}$ (or graph diffusion adjacency matrix $\mathbf{A}^{d}$) as input, which are essential for a deeper understanding of cellular diversity. This network requires to reconstruct both the weighted attribute matrix and the adjacency matrix simultaneously. Moreover, a layer in the encoder and decoder is formulated as:
$\mathbf{Z}_{C}^{(l)}=\sigma\left(\mathbf{A}^{r}\mathbf{Z}_{C}^{(l-1)}W_{CGA}^{(l)}\right),$ (16)
$\widehat{\mathbf{Z}}_{C}^{(h)}=\sigma\left(\mathbf{A}^{r}\widehat{\mathbf{Z}}_{C}^{(h-1)}\widehat{W}_{CGA}^{(h)}\right),$
where $W_{CGA}^{(l)}$ and $\widehat{W}_{CGA}^{(h)}$, respectively, denote the learnable parameters of $l$-th encoder layer and $h$-th decoder layer. $\sigma$ is a non-linear activation function, such as ReLU or Tanh.
To minimize both the reconstruction loss functions over GEA and CGA, our model is designed to minimize a hybrid loss function Tu et al. (2021):
[…]
where, $\eta$ is a pre-defined hyper-parameter. By minimizing Eq. 17, is termed to minimize the reconstruction loss over node attributes and graph structure adopted.
Denoising ZINB model-based Feature Decoder. The ZINB distribution models highly sparse and overdispersed count data Eraslan et al. (2019). To explain the characteristics of scRNA-seq data, we assume that the data follow ZINB distribution, namely:
[…]
$\small\operatorname{ZINB}(\mathbf{X}_{ij};\pi_{ij},\mu_{ij},\theta_{ij})=\pi_{ij}\delta_{0}(\mathbf{X}_{ij})+(1-\pi_{ij})\mathrm{NB}(\mathbf{X}_{ij};\mu,\theta),$ (19)
where dropout ($\pi_{ij}$) are the probability of dropout events, mean ($\mu_{ij}$) and dispersion ($\theta_{ij}$) parameters, respectively, represent the negative binomial component.
The decoder has three output layers to estimate parameters $\pi$, $\mu$ and $\theta$, which are
$\hat{\pi}=\operatorname{sigmoid}\left(W_{\pi}\cdot\mathbf{H}\right),$ (20)
$\hat{\mu}=diag\left(s_{i}\right)\times\exp\left(W_{\mu}\cdot\mathbf{H}\right),$
$\hat{\theta}=\exp\left(W_{\theta}\cdot\mathbf{H}\right).$
Similarly, $W_{\pi}$, $W_{\mu}$ and $W_{\theta}$ are the corresponding weights, the size factors $s_{i}$ is the ratio of the total cell count to the median $S$. Finally, we minimize the following overall negative likelihood loss of the ZINB model-based autoencoder, namely
$\mathcal{L}_{ZINB}=-log\left(ZINB\left(\mathbf{X}_{ij}\mid\hat{\pi}_{ij},\hat{\mu}_{ij},\hat{\theta}_{ij}\right)\right)$ (21)
C.4 Similarity Measurements
Embedded Points and Clustering Center. We use the Student’s t-distribution Van der Maaten and Hinton (2008) to measure the similarity between embedded points and the clustering center. The Student’s t-distribution Van der Maaten and Hinton (2008) is a popularly used kernel to measure the similarity between the embedded point $h_{i}$ and the clustering center $c_{j}$. It is defined as follows,
$q_{ij}=\frac{(1+\|h_{i}-c_{j}\|^{2}/\varepsilon)^{-\frac{1+\varepsilon}{2}}}{\sum_{j^{{}^{\prime}}}(1+\|h_{i}-c_{j^{{}^{\prime}}}\|^{2}/\varepsilon)^{-\frac{1+\varepsilon}{2}}},$ (22)
where $h_{i}=f(\bm{x}_{i})\in\mathbf{H}$ corresponds to $\bm{x}_{i}\in\widetilde{\mathbf{X}}$ after embedding, $\varepsilon$ is the degrees of freedom of the Student’s t-distribution, and $q_{ij}$ can be interpreted as the probability of assigning sample $i$ to cluster $j$ (i.e., a soft assignment). We define $Q=\left[q_{ij}\right]$ as the distribution of the assignments of all samples.
Data Representation and Clustering Center. We use the target distribution distance  Bo et al. (2020) to measure the similarity between data representation and the clustering center. Specifically, we calculated a target distribution $P=\left[p_{ij}\right]$:
$p_{ij}=\frac{q_{ij}^{2}/f_{j}}{\sum{q_{ij^{{}^{\prime}}}^{2}/f_{j^{{}^{\prime}}}}},$ (23)
where $f_{j}=\sum_{i}a_{ij}$ are soft cluster frequencies. In the target distribution $P$, each assignment in $Q$ is squared and normalized to enhance assignment confidence Bo et al. (2020).
Appendix D Experimental Details
D.1 Evaluation Metrics
Given the knowledge of the ground truth class assignments $U$ and our clustering algorithm assignment $V$ on $n$ data points.
Accuracy (ACC): The ACC is defined to measure the best matching between two cluster assignments $U$ and $V$. If there is a data point $i$, $l_{i}$ is the ground truth label, and $u_{i}$ is the clustering algorithm’s assignment, then the ACC is defined as:
$\text{ACC}=\max_{m}\frac{\sum_{i=1}^{n}1{l_{i}=m(u_{i})}}{n},$ (24)
here, $m$ ranges over all possible one-to-one mapping between $U$ and $V$, the best mapping with respect to $m$ can be efficiently found using the Hungarian algorithm Kuhn (1955).
Normalized Mutual Information (NMI): The NMI Strehl and Ghosh (2002) is a function that measures the consistency between the predicted and true labels of $n$ cells. Specifically,
$\small\text{NMI}=\frac{\sum_{i=1}^{C_{U}}\sum_{j=1}^{C_{V}}\frac{\mid U_{i}\bigcap V_{j}\mid}{n}\text{log}\frac{n\mid U_{i}\bigcap V_{j}\mid}{\mid U_{i}\mid\mid V_{j}\mid}}{\text{mean}(-\sum_{i=1}^{C_{U}}\mid U_{i}\mid\text{log}\frac{\mid U_{i}\mid}{n},-\sum_{j=1}^{C_{V}}\mid V_{j}\mid\text{log}\frac{\mid V_{j}\mid}{n})}.$ (25)
Adjusted Rand Index (ARI): The ARI Vinh et al. (2009) is a function that evaluates the similarity of the two assignments ignoring permutations. Given a set $X$ of size $n$ and two clustering results $U=\{u_{1},u_{2},\cdots,u_{r}\}$, $V=\{v_{1},v_{2},\cdots,v_{s}\}$, we denote $n_{ij}=\left|U_{i}\cap V_{j}\right|,i=1,2,\ldots,r,j=1,2,\cdots,s$. We also let $a_{i}=\sum_{i=1}^{r}n_{ij}$, $b_{j}=\sum_{j=1}^{s}n_{ij}$. Then we evaluate:
$\text{ARI}=\frac{\sum_{i,j}\binom{n_{i,j}}{2}-\left[\sum_{i}\binom{a_{i}}{2}\sum_{j}\binom{b_{j}}{2}\right]\binom{n}{2}}{\left[\sum_{i}\binom{a_{i}}{2}\sum_{j}\binom{b_{j}}{2}\right]/2-\left[\sum_{i}\binom{a_{i}}{2}\sum_{j}\binom{b_{j}}{2}\right]/\binom{n}{2}}.$ (26)
Macro Precision: It calculates the precision for each class (i.e., the proportion of correctly predicted positive instances for each class out of all instances predicted as that class), and then averages the precision over all classes.
Macro $\text{Macro Precision}=\frac{1}{N}\sum^{N}_{i=1}\frac{TP_{i}}{TP_{i}+FP_{i}}$ (27)
where $TP_{i}$ is the count of true positives for class $i$, $FP_{i}$ is the count of false positives for class $i$, and $N$ represents the total number of classes. This metric is particularly useful for assessing the performance of classification models, especially in multi-class scenarios.
Macro Recall: It calculates the recall for each class (i.e., the proportion of correctly predicted positive instances for each class out of all actual positive instances), and then averages the recall over all classes.
Macro $\text{Macro Recall}=\frac{1}{N}\sum^{N}_{i=1}\frac{TP_{i}}{TP_{i}+FN_{i}}$ (28)
where $TP_{i}$ is the number of true positives for the $i^{th}$ class, and $FN_{i}$ is the number of false negatives for the $i^{th}$ class. It indicates the model’s average ability to retrieve positive instances for each class.
Macro F1: It calculates the F1 score for each class (i.e., the harmonic mean of precision and recall for each class), and then averages the F1 scores over all classes.
[…]
It is a comprehensive metric used to evaluate the overall performance of a model, particularly in the presence of imbalanced classes.
D.2 Baseline Methods
Ten clustering competing methods are used in our experimental design. Among those approaches, pcaReduce Žurauskienė and Yau (2016) and SUSSC Wang et al. (2021a) stand out as early traditional clustering techniques. Additionally, we consider four deep neural network-based clustering methods, namely DEC Xie et al. (2016), scDeepCluster Tian et al. (2019), scDCC Tian et al. (2021), and scNAME Wan et al. (2022), which train an autoencoder to obtain representations followed by applying a clustering algorithm. scDSC Gan et al. (2022), scGNN Wang et al. (2021b), and scCDCG Xu et al. (2024) are representative GNN-based clustering methods that incorporate both node attributes and structural information to obtain latent representation for clustering.
  • Early Traditional Clustering Methods:
    • pcaReduce Žurauskienė and Yau (2016) merges PCA with K-means clustering and iteratively amalgamates clusters based on relevant probability density functions.
    • SUSSC Wang et al. (2021a) employs stochastic gradient descent together with a nearest neighbor search strategy to achieve efficient and accurate clustering.
  • Deep Neural Network-based Clustering Methods:
    • DEC Xie et al. (2016) is deep-embedded clustering, employing neural networks to learn both feature representations and cluster assignments.
    • scDeepCluster Tian et al. (2019), a single-cell model-based deep embedded clustering method, which simultaneously learns feature representation and clustering via explicit modeling of scRNA-seq data generation.
    • scDCC Tian et al. (2021) encodes prior domain knowledge as constraint information, integrating it effectively into the representation learning process using a novel loss function designed to enhance clustering accuracy.
    • scNAME Wan et al. (2022) integrates a mask estimation task alongside a neighborhood contrastive learning framework, optimizing both denoising and clustering procedures for scRNA-seq data.
  • Deep Structural Clustering Methods:
    • scDSC Gan et al. (2022) learns data representations using the ZINB model-based autoencoder and GNN modules, and uses mutual supervised strategy to unify these two architectures.
    • scGNN Wang et al. (2021b) utilizes GNNs and three multi-modal autoencoders to formulate and aggregate cell-cell relationships and models heterogeneous gene expression patterns using a left-truncated mixture Gaussian model.
    • scCDCG Xu et al. (2024) is a clustering framework for scRNA-seq data that combines cut-informed graph embedding, self-supervised learning with optimal transport, and autoencoder-driven feature learning to fully leverage complex intercellular relationships.
D.3 Classification Model
The classification model adopts a four-layer linear structure with ReLU activation functions and batch normalization between each layer. Specifically, we employed a Multilayer Perceptron (MLP) architecture consisting of input, two hidden, and output layers , ultimately predicting cell types. The dataset was meticulously partitioned into training and testing sets with an 8:2 ratio to ensure robust generalization during model training. Optimization objectives were guided by the CrossEntropyLoss function, and optimization was performed using the Adam optimizer with an initial learning rate set to 1e-2. Furthermore, we employed the ReduceLROnPlateau scheduler to dynamically adjust the learning rate based on loss minimization, utilizing a decay factor of 0.9 to expedite convergence to the optimal solution. Throughout the training phase, we meticulously conducted 100 epochs with validation checks performed every 10 epochs to meticulously monitor training progress and performance.
Appendix E Implementation Details
Training Procedure. The training process of scSiameseClu includes three steps: (i) we pre-train the GEA and CGA independently for at least 50 epochs by minimizing the reconstruction loss functions; (ii) both subnetworks are integrated into a united framework for another 100 epochs; (iii) Under the guidance of Eq.14 in the main text, we train the entire network for at least 50 epochs until convergence. As the number of cell clusters $k$ is unknown in real applications, we adopt the K-means method over the consensus clustering embedding $\mathbf{Z}$ to determine the optimal value of $k$. The code is available at the link: https://github.com/XPgogogo/scSiameseClu.
Parameters Setting. In practice, we implement scSiameseClu in Python 3.7 based on PyTorch. For competing methods, we reproduce their source code by following the setting of the original literature and present the average results. For our proposed method, we adopt DCRN Liu et al. (2022) as our backbone network. For all datasets, the trade-off parameter $\gamma$ was set to 1e3. The hyper-parameters $\beta$, $\lambda$, $\eta$, and $\varepsilon$ were initially set to 0.5, 0, 5, 0.1, and 1, respectively. Specific values were assigned to $\alpha$ for different datasets: 0.32 for CITE-CMBC, 0.25 for Human liver cells, 0.5 for Human kidney cells, and 0.1 for all other datasets. Similarly, $\rho$ was set to 100 for CITE-CMBC, 50 for Human kidney cells, and 10 for the remaining datasets. The value of $\sigma$ was adjusted to 5 for Human kidney cells, 1 for QS mouse lung cells, and 2 for the rest. The Adam optimizer was employed for optimization, with learning rates tailored to each dataset: 5e-3 for Human kidney cells, 5e-4 for QS mouse lung cells, and 1e-4 for all other datasets. To ensure the accuracy of the experimental results, we conducted each experiment 10 times with random seeds on all the datasets and reported the mean and variance of the results.
Appendix F Experiment Results
F.1 Additional Quantitative Analysis
By further analyzing the experimental results in Tab.1 in the main text, we can draw the conclusion: scSiameseClu significantly improves upon existing techniques, demonstrating superior representation learning capabilities on scRNA-seq data while successfully addressing the issue of representation collapse. Specifically, by thoroughly extracting intercellular high-order structural information and integrating gene expression attributes, scSiameseClu effectively learned discriminative features within the scRNA-seq data. Our approach improves data interpretability and robustly supports the discovery of complex biological relationships between cells, thereby facilitating further advancements in biological research.
(a) DEC
(b) scDeepCluster
(c) scDCC
(d) scCDCG
Figure 8: Dot plot of the top three DEGs for each cell type identified by four baseline models using the human pancreatic cells dataset.
F.2 Additional Biological Analysis.
(a) scDeepCluster
(b) scDCC
(c) scDSC
(d) scCDCG
(e) scSiameseClu
Figure 9: Cell type annotation: overlap of the top 50 DEGs in clusters from five methods versus gold standard cell types useing the human liver cells dataset. (similarity = overlapping DEGs/50).
Figure 10: Matrix plot of the top 3 DEGs for each cell type by  scSiameseCluuseing the human liver cells dataset.
Based on the previous experimental results, we have validated the capability of scSiameseClu to learn discriminative feature representations and achieve excellent clustering performance. Analyzing gene expression patterns through clustering of scRNA-seq data has proven advantageous in expediting the learning and application of various biological downstream tasks. Therefore, extracting meaningful biological insights from clustering results is essential for practical applications. Here, taking human pancreatic cells and Human liver cells as examples, we further conduct biological analysis to find out whether our method is beneficial for downstream analysis, including marker gene identification, cell type annotation, and cell type classification, based on the labels predicted by the scSiameseClu.
The gene expression matrix and predicted labels were utilized to identify DEGs and further recognize marker genes for each cluster. Standard cell type annotation methods were applied to Human pancreatic cells and Human liver cells datasets, assigning cell types to each cluster based on established markers and cluster-specific gene expression profiles. We employed the "FindAllMarkers" function in the Seurat Butler et al. (2018) package with default parameters to identify DEGs for each cluster, thereby obtaining DEGs within the gold standard clusters annotated with respective cell types. Differentially expressed genes (DEGs) are important indicators for the cell type annotations of the clusters obtained through different methods. To determine whether scSiameseClu can accurately annotate clusters, we processed scSiameseClu and each competing method identically, identifying the top 50 marker genes for each cluster and comparing their overlap with the gold standard, referred to as the similarity of DEGs between different methods and the gold standard (as shown in Fig.5 in the main text, scSiameseClu could assign a cell type to a specific cluster, with most clusters showing a differential gene similarity greater than 90%. For example, cluster 1 was identified as endothelial cells, and cluster 2 as type B pancreatic cells, etc. scDeepCluster could not identify the cell type for cluster 1; scNAME could not identify cell types for clusters 2 and 7, nor could it assign a unique cell type to cluster 0; scGNN and scCDCG could not assign endothelial cell to a specific cluster.
Detailed results showed that other methods, such as scDeepCluster, scNAME, scGNN, and scCDCG, failed to consistently identify specific cell types for various clusters, highlighting the superiority of our approach in accurately matching clusters with known cell types. To further DEGs identified by the scSiameseClu clustering, Fig.6 in the main text depicts a dot plot of the top 3 DEGs for each of the 9 clusters. These genes represent the most characteristic DEGs within each cluster identified by our method and can be considered as the marker genes for the respective clusters.
In further analysis of the human pancreatic cells dataset, we plotted the top three differentially expressed genes (DEGs) for each cell type from four baseline models (DEC, scDeepCluster, scDCC, scSDCG), as shown in Fig. 8. Comparing these results with Fig.6 in the main text, we found that scSiameseClu excels in identifying cell type-specific DEGs, particularly demonstrating greater consistency and reliability in the selection of highly expressed genes. This provides clear support for subsequent biological research. Fig. 9 and Fig. 10 in the supplementary materials are supplemented with experimental results conducted on Human liver cells dataset.
Therefore, our clustering results can help researchers identify marker genes and annotate cells, which is able to further facilitate the understanding of different cell types, uncover biological differences between cells, and provide crucial insights for downstream functional analysis and mechanistic studies.
F.3 Computational Efficiency
The running time of scSiameseClu and seven baselines (DEC, scDeepCluster, scDCC, scNAME, scDSC, scGNN, and scCDCG). The results show that scSiameseClu consistently demonstrates superior computational efficiency across various datasets, while baseline methods experience varying computational burdens. The improved efficiency of scSiameseClu reduces training time and computational resource requirements, making it a viable and powerful tool for practical applications in biological data analysis.
Figure 11: The running time of scSiameseClu and seven baselines.
F.4 Scalability of scSiameseClu
With the increasing number of cells being profiled in scRNA-seq experiments, there is an urgent need to develop analytical methods that can effectively handle large datasets. To assess the scalability of scSiameseClu, we summarized the runtime of  scSiameseClu on simulation datasets of varying sizes. We designed the following simulations under extensive settings approximating different biological scenarios. Specifically, we applied the R package Splatter Zappia et al. (2017) to simulate seven simulated datasets, repsectively, containing 2k, 4k, 8k, 16k, 32k cells with 3k genes. Fig. 12 reports the runtime for one epoch of scSiameseClu at different stages (i.e., AE, GAE, pre-train, and fine-tuning) on these datasets. It can be observed that as the number of cells increases, the runtime increases almost linearly, with the most significant increase occurring during the fine-tuning stage.
Figure 12: Applying scSiameseClu on various simulated data.
(a) $\alpha$
(b) $\beta$
(c) $\gamma$
(d) $\lambda$
(e) $\eta$
(f) $\sigma$
Figure 13: Performance of scSiameseClu on Human Pancreas cell benchmark datasets w.r.t. different hyper-parameters, including $\alpha$, $\beta$, $\gamma$, $\lambda$, $\eta$, and $\sigma$.
F.5 Parameter sensitivity
Our work introduces seven hyperparameters: $\alpha$, $\beta$, $\gamma$, $\lambda$, $\mu$, $\delta$, and $\sigma$, which balance the contributions of different modules. Here are the details:
  • $\alpha$: Probability of message passing through higher-order information for generating the graph diffusion matrix.
  • $\beta$: Proportion parameter that controls the fusion ratio of $\mathbf{Z}^{E}$ and $\mathbf{Z}^{R}$; its weight is learned during training.
  • $\lambda$: Smoothing parameter controlling cluster balance in the optimal transport strategy.
  • $\eta$, $\delta$, and $\sigma$: Parameters controlling the contributions of different modules in the loss function.
As shown in the Fig. 13, we examined the impact of these hyperparameters on model performance through a sensitivity analysis using Human Pancreas cell dataset. These results indicate that scSiameseClu maintains stable performance across a wide range of hyperparameter values. For instance, values of between $0.1$ and $0.9$ all yielded ACC values around $0.95$, demonstrating robustness to this hyperparameter. Similarly, the other hyperparameters also show stable performance, which is critical for practical applications where tuning hyperparameters can be challenging. To enhance the reproducibility of experiments and the generalizability of the model, we recommend initially setting a group of hyperparameters: $\alpha=0.1$, $\beta=1$, $\gamma=1e3$, $\lambda=5$, $\eta=10$, and $\sigma=2$. These values serve as starting points for hyperparameter tuning, and adjusting them based on the specific characteristics of the data is necessary when applying scSiameseClu to new datasets. Although our sensitivity analysis indicates that these settings generally yield near-optimal results across various datasets, fine-tuning the parameters to accommodate different datasets helps ensure the model’s stability and optimal performance.
Appendix G Related Work
G.1 Traditional Clustering Algorithm
In the realm of scRNA-seq data analysis, early traditional clustering algorithms typically employ hard clustering methods, which involve dimensionality reduction followed by clustering using algorithms like k-means or hierarchical clustering. Within these traditional approaches, Phenograph Levine et al. (2015), MAGIC Van Dijk et al. (2018), and Seurat Butler et al. (2018) utilize k-nearest neighbor (KNN) graphs to simulate relationships between cells, with MAGIC being the first method to explicitly impute the entire genome in single-cell gene expression profiles. In contrast, CIDR Lin et al. (2017) adopts an implicit imputation method to mitigate the impact of dropout events. SIMLR Wang et al. (2018) and MPSSC Park and Zhao (2018), on the other hand, are multi-kernel spectral clustering methods that leverage multiple kernel functions to learn robust similarity measures corresponding to different information representations in the data. MPSSC is well suited to represent the sparsity of scRNA-seq data. Additionally, RaceID Lin et al. (2017) introduces outlier detection to enhance the ability of K-means to identify subpopulation cells, while scImpute Li and Li (2018) is a statistical method that aims to accurately and robustly impute dropout values in scRNA-seq data. However, these methods often rely on hard clustering algorithms, making it difficult to handle high-dimension data and capture complex nonlinear data features, thus making them susceptible to inherent noise in the data.
G.2 Deep Feature Clustering Algorithm
Deep learning has also gained prominence in scRNA-seq clustering. Methods like scScope Deng et al. (2018) and DCA Eraslan et al. (2019) are based on autoencoders, obtaining effective feature representations through supervised learning. DCA introduces an autoencoder based on the zero-inflated negative binomial (ZINB) model to simulate the distribution of scRNA-seq data, generating a representation matrix to capture nonlinear dependencies between genes. SCVI Lopez et al. (2018) and SCVIS Ding et al. (2018), which rely on autoencoders for dimensionality reduction, often face over-regularization due to their Gaussian distribution assumption. In contrast, DeepImpute Arisdakessian et al. (2019), based on standard deep neural networks, focuses on imputing missing values. DEC Xie et al. (2016) and DESC Li et al. (2020) employ neural networks to learn feature representations and clustering assignments. DESC gradually eliminates batch effects through self-learning iterations, reducing technical differences between batches to approach real biological variations. However, these methods primarily focus on denoising and imputing dropout values in scRNA-seq data. Additionally, scDeepCluster Tian et al. (2019) combines ZINB-based autoencoders with deep embedding clustering to optimize latent feature learning and clustering for better results. scDCC Tian et al. (2021) encodes prior knowledge as constraint information and integrates it into the representation learning process through a novel loss function. AttentionAE-sc Li et al. (2023) combines denoising representation learning with cluster-friendly representation learning using attention mechanisms. However, the aforementioned approaches predominantly concentrate on extracting features from isolated cells, neglecting the vital intercellular structural information, crucial for accurately representing the variations between different cell types.
G.3 Deep Structural Clustering Algorithm
Recently, deep structural clustering methods have gained considerable attention and achieved remarkable effectiveness. scGAE Luo et al. (2021) and graph-sc Ciortan and Defrance (2022) utilize graph autoencoders to embed scRNA-seq data while preserving their topological structures. scVGAE Inoue (2024) integrates GNNs into a variational autoencoder framework and utilizes a ZINB loss function. scDeepSort Shao et al. (2021) pioneers the use of pre-trained GNN models for annotating cell types in scRNA-seq data and introduces a strategy using gene expression values to construct cell-gene graphs. scGAC Cheng and Ma (2022) constructs cell graphs and adopts a self-optimization approach for simultaneous learning of representations and clustering optimization. Notably, scGNN Wang et al. (2021b) utilizes GNNs and multi-modal autoencoders to aggregate cell-cell relationships and model gene expression patterns using a left truncated mixture Gaussian model, scDSC Gan et al. (2022) combines a ZINB model-based autoencoder with GNN modules, employing a mutually supervised strategy for enhanced data representation. GraphSCC Zeng et al. (2020) employs GNNs to capture structural cell-cell relationships, optimizing the network’s output with a self-supervised module. However, these GNN-based methods often face issues like representation collapse, and their high complexity and limited scalability often hinder their application to large-scale scRNA-seq datasets.